# The Non-Gaussian to Gaussian Transition: Pointwise Heat Kernel Estimates and Optimal Convergence Rates

### 15 Mar 2026 v1

@@ 2.1.2. Infinitesimal generators and Lévy measures
It is well-known that the infinitesimal generator of a standard Brownian motion is the Laplacian $\Delta:=\sum_{i=1}^{d}\partial_{x_{i}}^{2}$, and the one of $\alpha$-stable process is the fractional Laplacian operator $\Delta^{\alpha/2}:=-(-\Delta)^{\alpha/2}$ defined by
@@
We conclude this subsubsection by presenting some trivial but useful facts about Lévy measure $\nu^{(\alpha)}$ for $\alpha\in(0,2)$. Since $\lim_{s\rightarrow 0}s\Gamma(s)=\lim_{s\rightarrow 0}\Gamma(s+1)=1$, we obtain that
$\lim_{\alpha\rightarrow 2}\frac{\omega_{d-1}}{d}\frac{\mathcal{C}(d,\alpha)}{2(2-\alpha)}=1,$
where $\mathcal{C}(d,\alpha)$ is defined by (1.12). Furthermore, by Lemma 2.2 of  [DSX23], there is a constant $c>0$ independent of $\alpha$ such that
@@
$\left|\frac{\omega_{d-1}\mathcal{C}(d,\alpha)}{d(2-\alpha)}-1\right|\leq c(2-\alpha)\log(1+d).$ (2.12)
Hence, fixed $\delta>0$, for any $\vartheta>\alpha$,
@@ 2.2. Some basic inequalities of $\varrho^{(k)}_{\gamma_{1},\gamma_{2}}(t,x)$
Assume that $j=0,1$ and $\alpha\in[\alpha_{0},2)$ with some $\alpha_{0}\in(j,2)$. Then there is a constant $c=c(d,\alpha_{0})>0$ such that for $s>0$ and $x\in{\mathbb{R}}^{d}$,
$\int_{|z|>s^{1/\alpha}\vee\frac{|x|}{2}}(s^{-\frac{1}{\alpha}}|z|)^{j}\varrho_{\alpha}(s,x+z)\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)\leq c(2-\alpha)s^{-1}\left(1+s^{-j/\alpha}|x|^{j}\right)\varrho_{\alpha}(s,x).$ (2.24)
@@
${\lesssim}(2-\alpha)s^{-j/\alpha}|x|^{-d-\alpha}\int_{{\mathbb{R}}^{d}}\left(|x+z|^{j}+|x|^{j}\right)\varrho_{\alpha}(s,x+z){\mathord{{\mathrm{d}}}}z$
${\lesssim}(2-\alpha)s^{-j/\alpha}|x|^{-d-\alpha}\left(\int_{{\mathbb{R}}^{d}}|z|^{j}\varrho_{\alpha}(s,z){\mathord{{\mathrm{d}}}}z+\int_{{\mathbb{R}}^{d}}|x|^{j}\varrho_{\alpha}(s,z){\mathord{{\mathrm{d}}}}z\right)$
[…]
@@ 2.3. Some uniform estimates of densities
Assume that $\alpha\in[\alpha_{0},2]$ with some $\alpha_{0}\in(0,2)$. There is a constant $c>0$ only depends on $d,\alpha_{0}$ such that, for any $t\geq 0$,
@@
Indeed, recalling the elementary fact that for any real number $n>0$,
$\mathrm{e}^{-u}\leq n^{n}u^{-n},\ \ \forall u\geq 0,$
one sees that for any real number $n\geq 0$ and constant $c>0$,
$\mathrm{e}^{-\frac{|x|^{2}}{c}}\leq(cn)^{n}|x|^{-2n},$ (2.28)
where we used the conventions $0^{0}=1$ and $\frac{1}{0}=\infty$. From (2.28), picking $c=4t$, and taking $n=0$ and $n=(d+2)/2$ in turn, we get the desired estimate (2.27) by (2.3).
@@ 3. Heat kernels: uniform estimates and $\alpha$-continuity
$\widetilde{{\mathscr{L}}}_{r}^{(\alpha),\sigma}f(x):=\begin{cases}\int_{{\mathbb{R}}^{d}}{\Big(}f(x+\sigma(r)z)-f(x)-\sigma(r)z{\boldsymbol{1}}_{|z|\leq 1}\cdot\nabla f(x){\Big)}\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z),&\ \ \text{when}\ \ \alpha\in(0,2),\\ \frac{1}{2}{\rm TR}{\Big(}(\sigma\sigma^{*})(r){\rm H}(f)(x){\Big)},&\ \ \text{when}\ \ \alpha=2,\end{cases}$ (3.4)
@@ 3.1. Uniform heat kernel estimates for non-local and local
Assume that $\alpha\in[\alpha_{0},2]$ with some $\alpha_{0}\in(0,2)$, and $j\in{\mathbb{N}}_{0}$. For any $0\leq s<\infty$ and $t>0$, there is a constant $c>0$ only depends on $d,j,\alpha_{0},\kappa_{1}$ such that
@@
$|\nabla^{j}p^{(2),\sigma}(t,x)-\nabla^{j}p^{(2),\sigma}(t,\bar{x})|\leq c((t^{-1/2}|x-\bar{x}|)\wedge 1){\Big(}p^{(2)}(4t,x)+p^{(2)}(4t,\bar{x}){\Big)}.$
@@ 3.2. $\alpha$-stability estimates of infinitesimal generators
Using the definition of operators, we have that for any $\epsilon>0$ (taking $\epsilon=1$ for simplicity),
@@
$I_{1}=$ $\int_{|z|\leq\epsilon}\left[\int_{0}^{1}{\mathord{{\mathrm{d}}}}r\int_{0}^{r}\nabla^{2}f(x+s\sigma(t)z)(\sigma(t)z)\cdot\sigma(t)z{\mathord{{\mathrm{d}}}}s\right]\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)\widetilde{{\mathscr{L}}}^{(2),\sigma}_{t}f(x)$
@@
$+\int_{|z|\leq\epsilon}\int_{0}^{1}{\mathord{{\mathrm{d}}}}r\int_{0}^{r}\left[\nabla^{2}f(x+s\sigma(t)z)-\nabla^{2}f(x)\right](\sigma(t)z)\cdot\sigma(t)z{\mathord{{\mathrm{d}}}}s\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)$
@@
For $I_{11}$, using (2.12), we obtain that
@@
The following result is a stronger version of Lemma 3.4, and will be used in the proofs of Corollary 3.6 and Corollary 4.5.
@@
Assume that $\sigma_{1},\sigma_{2}$ satisfy the condition (3.1). If $\alpha\in[\alpha_{0},2)$ with some $\alpha_{0}\in(1,2)$, then there is a constant $c=c(d,\kappa_{1},\alpha_{0})>0$ such that for $t>r>s$,
$|(\widetilde{\mathscr{D}}^{(\alpha),\sigma_{1}}_{r}-\widetilde{\mathscr{D}}^{(\alpha),\sigma_{2}}_{r})$ $p^{(\alpha),\sigma}|(s,t,y)\leq c(2-\alpha)\left[1\wedge\left(\|\sigma_{1}(r)-\sigma_{2}(r)\|(1+(t-s)^{-1/\alpha}|y|)\right)\right]$
$\times(1+|\ln(t-s)|)(|t-s|^{-1}+|t-s|^{-\frac{2}{\alpha}})\varrho_{\alpha}(t-s,y),$
and
$|\widetilde{\mathscr{D}}^{(\alpha),\sigma}_{r}p^{(\alpha),\sigma}|(s,t,y)$ $\leq c(2-\alpha)(1+|\ln(t-s)|)(|t-s|^{-1}+|t-s|^{-\frac{2}{\alpha}})\varrho_{\alpha}(t-s,y),$
where $\widetilde{\mathscr{D}}^{(\alpha),\sigma}_{r}$ is defined by (3.16).
@@
Since the second one is similar and easier, we only prove the first one. For the simplicity of notations, we drop the time variable $r$. Define the first-order difference operator $\delta_{z}$ by
@@
First of all, by (3.1), we have
$\Big|\big(\delta_{\sigma_{1}z}\nabla^{2}f(y)\big)(\sigma_{1}z)\cdot\sigma_{1}z$ $-\big(\delta_{\sigma_{2}z}\nabla^{2}f(y)\big)(\sigma_{2}z)\cdot\sigma_{2}z\Big|\lesssim|\delta_{\sigma_{1}z}\nabla^{2}f(y)-\delta_{\sigma_{2}z}\nabla^{2}f(y)||z|^{2}$
$+\left|{\Big(}\delta_{\sigma_{2}z}\nabla^{2}f(y){\Big)}(\sigma_{1}z)\cdot\sigma_{1}z-{\Big(}\delta_{\sigma_{2}z}\nabla^{2}f(y){\Big)}(\sigma_{2}z)\cdot\sigma_{2}z\right|.$ (3.17)
Following the proof of Lemma 3.4, a similar argument derives that for $i=1,2$,
$\widetilde{\mathscr{D}}^{(\alpha),\sigma_{i}}f(x)=$ $\left(\frac{1}{d}\int_{|z|\leq\varepsilon}\left[\int_{0}^{1}{\mathord{{\mathrm{d}}}}\theta\int_{0}^{\theta}\widetilde{{\mathscr{L}}}^{(2),\sigma_{i}}f(x)|z|^{2}{\mathord{{\mathrm{d}}}}\vartheta\right]\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)-\widetilde{\mathscr{L}}^{(2),\sigma_{i}}f(x)\right)$
$+\int_{|z|\leq\varepsilon}\int_{0}^{1}{\mathord{{\mathrm{d}}}}\theta\int_{0}^{\theta}\left[\delta_{\vartheta\sigma_{i}z}\nabla^{2}f(x)\right](\sigma_{i}z)\cdot\sigma_{i}z{\mathord{{\mathrm{d}}}}\vartheta\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)$
$+\int_{|z|>\varepsilon}{\Big[}\delta_{\sigma_{i}z}f(x){\Big]}\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z).$
and then, taking $\varepsilon=|t-s|^{1/\alpha}>0$, by (3.2), one sees that
$|($ $\widetilde{\mathscr{D}}^{(\alpha),\sigma_{1}}-\widetilde{\mathscr{D}}^{(\alpha),\sigma_{2}})p^{(\alpha),\sigma}_{s,t}(y)|\lesssim\xi^{(\alpha)}_{s,t}\times|(\widetilde{{\mathscr{L}}}^{(2),\sigma_{1}}-\widetilde{{\mathscr{L}}}^{(2),\sigma_{2}})p^{(\alpha),\sigma}_{s,t}(y)|$
$+\int_{|z|\leq\varepsilon}\int_{0}^{1}{\mathord{{\mathrm{d}}}}\theta\int_{0}^{\theta}\left|\delta_{\vartheta\sigma_{1}z}\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y)-\delta_{\vartheta\sigma_{2}z}\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y)\right|{\mathord{{\mathrm{d}}}}\vartheta|z|^{2}\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)$
$+\int_{|z|\leq\varepsilon}\int_{0}^{1}{\mathord{{\mathrm{d}}}}\theta\int_{0}^{\theta}\left|\left[\delta_{\vartheta\sigma_{1}z}\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y)\right](\sigma_{1}z)\cdot\sigma_{1}z-\left[\delta_{\vartheta\sigma_{2}z}\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y)\right](\sigma_{2}z)\cdot\sigma_{2}z\right|{\mathord{{\mathrm{d}}}}\vartheta\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)$
$+\int_{|z|>\varepsilon}\left|p^{(\alpha),\sigma}_{s,t}(y+\sigma_{1}z)-p^{(\alpha),\sigma}_{s,t}(y+\sigma_{2}z)\right|\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z),$
@@
$\xi_{s,t}^{(\alpha)}$ $:=(t-s)^{\frac{2-\alpha}{\alpha}}\frac{\omega_{d-1}\mathcal{C}(d,\alpha)}{2d(2-\alpha)}-1$
@@
(Step 1) Observe that, when $|z|\leq\varepsilon$, under the condition (3.1), by (3.8) and (2.21), we have
$\left|\delta_{\vartheta\sigma_{1}z}\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y)-\delta_{\vartheta\sigma_{2}z}\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y)\right|$ $\leq\left|\delta_{\vartheta\sigma_{1}z}\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y)\right|+\left|\delta_{\vartheta\sigma_{2}z}\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y)\right|$
$\lesssim|t-s|^{-1/\alpha}|z|\varrho^{(2)}_{\alpha}(t-s,y)$
and
$\left|\delta_{\vartheta\sigma_{1}z}\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y)-\delta_{\vartheta\sigma_{2}z}\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y)\right|$ $=\left|\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y+\vartheta\sigma_{1}z)-\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y+\vartheta\sigma_{2}z)\right|$
$\lesssim|t-s|^{-1/\alpha}|(\sigma_{1}-\sigma_{2})z|\varrho^{(2)}_{\alpha}(t-s,y)$
$\leq\|\sigma_{1}-\sigma_{2}\|(t-s)^{-1/\alpha}|z|\varrho^{(2)}_{\alpha}(t-s,y),$

### 26 Mar 2026 v2

@@ 3.1.2. Infinitesimal generators and Lévy measures
It is well-known that the infinitesimal generator of a Brownian motion is the Laplacian $\Delta:=\sum_{i=1}^{d}\partial_{x_{i}}^{2}$, and the one of $\alpha$-stable process is the fractional Laplacian operator $\Delta^{\alpha/2}:=-(-\Delta)^{\alpha/2}$ defined by
@@
We conclude this subsubsection by presenting some trivial but useful facts about Lévy measure $\nu^{(\alpha)}$ for $\alpha\in(0,2)$. By Lemma 2.2 of  [DSX23], there is a constant $c>0$ independent of $\alpha\in(0,2)$ such that
@@
$\left|\frac{\omega_{d-1}\mathcal{C}(d,\alpha)}{d(2-\alpha)}-1\right|\leq c(2-\alpha)\log(1+d),$ (3.12)
where $\mathcal{C}(d,\alpha)$ is defined by (1.9). Hence, fixed $\delta>0$, for any $\vartheta>\alpha$,
@@ 3.2. Some basic inequalities of $\varrho^{(k)}_{\gamma_{1},\gamma_{2}}(t,x)$
Assume that $j=0,1$ and $\alpha\in[\alpha_{0},2)$ with some $\alpha_{0}\in(0,2)$. Then there is a constant $c=c(d,\alpha_{0})>0$ such that for $s>0$ and $x\in{\mathbb{R}}^{d}$,
$\int_{|z|>s^{1/\alpha}\vee\frac{|x|}{2}}\varrho_{\alpha}(s,x+z)\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)\leq c(2-\alpha)s^{-1}\varrho_{\alpha}(s,x).$ (3.24)
@@
$\lesssim(2-\alpha)|x|^{-d-\alpha}\int_{{\mathbb{R}}^{d}}\varrho_{\alpha}(s,x+z){\mathord{{\mathrm{d}}}}z$
@@ 3.3. Some uniform estimates of densities
First of all, recalling the elementary fact that for any real number $n>0$,
$\mathrm{e}^{-u}\leq n^{n}u^{-n},\ \ \forall u\geq 0,$
one sees that for any real number $n\geq 0$ and constant $c>0$,
$\mathrm{e}^{-\frac{|x|^{2}}{c}}\leq(cn)^{n}|x|^{-2n},$ (3.26)
where we used the conventions $0^{0}=1$ and $\frac{1}{0}=\infty$.
@@
Assume that $\alpha\in[\alpha_{0},2]$ with some $\alpha_{0}\in(0,2)$. There is a constant $c>0$ depending only on $d,\alpha_{0}$ such that, for any $t>0$,
@@
Indeed, from (3.26), picking $c=4t$, and taking $n=0$ and $n=(d+2)/2$ in turn, we get the desired estimate (3.28) by (3.3).
@@
We also need the following lemma, which will be applied to prove Corollary 4.4.
Lemma 3.7.
Assume that $T>0$ and $\alpha\in[\alpha_{0},2]$ with some $\alpha_{0}\in(0,2)$. There is a constant $c>0$ depending only on $d,T,\alpha_{0}$ such that, for any $t\in(0,T]$,
$p^{(2)}(t,x)\leq ct^{\frac{\alpha-2}{2}}\varrho_{\alpha}(t,x).$
Proof.
From (3.26), picking $c=4t$, and taking $n=0$ and $n=(d+\alpha)/2$ in turn, we get that
$p^{(2)}(t,x)$ $=(4\pi t)^{-\frac{d}{2}}\mathrm{e}^{-\frac{|x|^{2}}{4t}}\leq(4\pi t)^{-\frac{d}{2}}{\Big(}1\wedge{\Big[}(4t\tfrac{d+\alpha}{2})^{\frac{d+\alpha}{2}}|x|^{-d-\alpha}{\Big]}{\Big)}\lesssim t^{-\frac{d}{2}}\wedge\frac{t^{\alpha/2}}{|x|^{d+\alpha}},$
which implies that
$p^{(2)}(t,x)\lesssim(t^{\frac{d}{\alpha}-\frac{d}{2}}+t^{\frac{\alpha-2}{2}})\varrho_{\alpha}(t,x)\lesssim t^{\frac{\alpha-2}{2}}\varrho_{\alpha}(t,x).$
The proof is finished. ∎
@@ 4. Heat kernels: uniform estimates and $\alpha$-continuity
$\quad\widetilde{{\mathscr{L}}}_{r}^{(\alpha),\sigma}f(x)$
$:=\begin{cases}\int_{{\mathbb{R}}^{d}}{\Big(}f(x+\sigma(r)z)-f(x)-\sigma(r)z{\boldsymbol{1}}_{|z|\leq 1}\cdot\nabla f(x){\Big)}\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z),&\ \ \text{when}\ \ \alpha\in(0,2),\\ {\rm TR}{\Big(}(\sigma\sigma^{*})(r){\rm H}(f)(x){\Big)},&\ \ \text{when}\ \ \alpha=2,\end{cases}$ (4.4)
@@ 4.1. Uniform heat kernel estimates for non-local and local
Assume that $\alpha\in[\alpha_{0},2]$ with some $\alpha_{0}\in(0,2)$, and $j\in{\mathbb{N}}_{0}$. For any $0\leq s<\infty$ and $t>0$, there is a constant $c>0$ depending only on $d,j,\alpha_{0},\kappa_{1}$ such that
@@
$|\nabla^{j}p^{(2),\sigma}(t,x)-\nabla^{j}p^{(2),\sigma}(t,\bar{x})|\leq ct^{-j/2}{\Big(}(t^{-1/2}|x-\bar{x}|)\wedge 1{\Big)}{\Big(}p^{(2)}(4t,x)+p^{(2)}(4t,\bar{x}){\Big)}.$
@@
From Lemmas 4.3 and 3.7, we directly have the following corollary, which will be used in the proof of Lemma 4.6.
Corollary 4.4.
Assume that $T>0$ and $\alpha\in[\alpha_{0},2]$ with some $\alpha_{0}\in(0,2)$. For $j\in{\mathbb{N}}_{0}$, there is a constant $c>0$ depending only on $d,T,\alpha_{0},j,\kappa_{1}$ such that, for any $t\in(0,T]$,
$|\nabla^{j}p^{(2),\sigma}(t,x)|\leq c(t^{-1/2}\wedge|x|^{-1})^{j}t^{\frac{\alpha-2}{\alpha}}\varrho_{\alpha}(t,x),$
and
$|\nabla^{j}p^{(2),\sigma}(t,x)-\nabla^{j}p^{(2),\sigma}(t,\bar{x})|\leq c{\Big(}(t^{-1/2}|x-\bar{x}|)\wedge 1{\Big)}t^{\frac{\alpha-2}{\alpha}-\frac{j}{2}}{\Big(}\varrho_{\alpha}(t,x)+\varrho_{\alpha}(t,\bar{x}){\Big)}.$
@@ 4.2. $\alpha$-stability estimates of infinitesimal generators
Using the definition of operators, we have that for any $\epsilon>0$,
@@
$I_{1}=$ $\int_{|z|\leq\epsilon}\left[\int_{0}^{1}{\mathord{{\mathrm{d}}}}r\int_{0}^{r}\nabla^{2}f(x+s\sigma(t)z)(\sigma(t)z)\cdot\sigma(t)z{\mathord{{\mathrm{d}}}}s\right]\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)-\widetilde{{\mathscr{L}}}^{(2),\sigma}_{t}f(x)$
@@
$+\int_{|z|\leq\epsilon}\left[\int_{0}^{1}{\mathord{{\mathrm{d}}}}r\int_{0}^{r}\left[\nabla^{2}f(x+s\sigma(t)z)-\nabla^{2}f(x)\right](\sigma(t)z)\cdot\sigma(t)z{\mathord{{\mathrm{d}}}}s\right]\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)$
@@
In the following, take $\epsilon=1$ for simplicity. For $I_{11}$, using (3.13) and (3.12), we obtain that
@@
The following result will be used in the proofs of Corollary 4.8 and Corollary 5.5.
@@
Assume that $T>0$ and $\sigma_{1},\sigma$ satisfy the condition (4.1). If $\alpha\in[\alpha_{0},2)$ with some $\alpha_{0}\in(0,2)$, then there is a constant $c=c(d,T,\kappa_{1},\alpha_{0})>0$ such that for $0\leq s<r<t\leq T$,
$|\widetilde{\mathscr{D}}^{(\alpha),\sigma_{1}}_{r}p^{(\gamma),\sigma}|(s,t,y)$ $\leq c(2-\alpha)(1+|\ln(t-s)|)(|t-s|^{-1}+|t-s|^{-\frac{2}{\alpha}})\varrho_{\alpha}(t-s,y),$
where $\gamma\in\{\alpha,2\}$, and $\widetilde{\mathscr{D}}^{(\alpha),\sigma}_{r}$ is defined by (4.16).
@@
For the simplicity of notations, we drop the time variable $r$. Define the first-order difference operator $\delta_{z}$ by
@@
Following the proof of Lemma 4.5, a similar argument derives that
$\widetilde{\mathscr{D}}^{(\alpha),\sigma_{1}}f(y)=$ $\left(\frac{1}{d}\int_{|z|\leq\varepsilon}\left[\int_{0}^{1}{\mathord{{\mathrm{d}}}}\theta\int_{0}^{\theta}\widetilde{{\mathscr{L}}}^{(2),\sigma_{1}}f(y)|z|^{2}{\mathord{{\mathrm{d}}}}\vartheta\right]\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)-\widetilde{\mathscr{L}}^{(2),\sigma_{1}}f(y)\right)$
$+\int_{|z|\leq\varepsilon}\int_{0}^{1}{\mathord{{\mathrm{d}}}}\theta\int_{0}^{\theta}\left[\delta_{\vartheta\sigma_{1}z}\nabla^{2}f(y)\right](\sigma_{1}z)\cdot\sigma_{1}z{\mathord{{\mathrm{d}}}}\vartheta\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)$
$+\int_{|z|>\varepsilon}{\Big[}\delta_{\sigma_{1}z}f(y){\Big]}\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)$
In the following, we take $\varepsilon=|t-s|^{1/\alpha}>0$, and estimate these three terms in turn.
(Part 1) We prove the result for the case of $f(y)=p^{(\alpha),\sigma}_{s,t}(y)$ in this part.
(i) Under the condition (4.1), one sees that
[…]
@@
$\xi_{s,t}^{(\alpha)}$ $:=\left|(t-s)^{\frac{2-\alpha}{\alpha}}\frac{\omega_{d-1}\mathcal{C}(d,\alpha)}{d(2-\alpha)}-1\right|$
@@
(ii) Observe that, when $|z|\leq|t-s|^{1/\alpha}$, under the condition (4.1), we have
[…]
Thus, we have
$|\Lambda^{(2)}_{\alpha}(s,t,y)|$ $\lesssim|t-s|^{-3/\alpha}\varrho_{\alpha}(t-s,y)\left(\int_{|z|\leq\varepsilon}|z|^{3}\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)\right)$
[…]
(iii) Under the condition (4.1), by (4.8), we have
$|\delta_{\sigma_{1}z}p^{(\alpha),\sigma}_{s,t}(y)|\lesssim\varrho_{\alpha}(t-s,y+\sigma_{1}z)+\varrho_{\alpha}(t-s,y).$
Thus, we obtain that
$\Lambda^{(3)}_{\alpha}(s,t,y$ $)\leq\left(\int_{|z|>|t-s|^{1/\alpha}\vee\frac{|y|}{2\kappa_{1}}}+\int_{\frac{|y|}{2\kappa_{1}}\geq|z|>|t-s|^{1/\alpha}}\right)\varrho_{\alpha}(t-s,y+\sigma_{1}z)\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)$
[…]
$\qquad+2\varrho_{\alpha}(t-s,y)\int_{|z|>|t-s|^{1/\alpha}}\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)$
$\,\,\lesssim(2-\alpha)|t-s|^{-1}\varrho_{\alpha}(t-s,y),$
where we used Lemma 3.3 and (3.14) in the last inequality. Combining the above calculations and observing (4.17), the case of $\gamma=\alpha$ follows.
(Part II) In this part, we give the proof of the case of $f(y)=p^{(2),\sigma}_{s,t}(y)$. The proof is similar to that in Part I.
(i) Under the condition (4.1), by Corollary 4.4, one sees that
$|\widetilde{{\mathscr{L}}}^{(2),\sigma_{1}}p^{(2),\sigma}_{s,t}(y)|\lesssim|t-s|^{-1}|t-s|^{\frac{\alpha-2}{\alpha}}\varrho_{\alpha}(t-s,x)=|t-s|^{-\frac{2}{\alpha}}\varrho_{\alpha}(t-s,x),$
which derives that
$|\Lambda^{(1)}_{\alpha}(s,t,y)|\lesssim\xi_{s,t}^{(\alpha)}|t-s|^{-\frac{2}{\alpha}}\varrho_{\alpha}(t-s,x).$
(ii) Observe that, by the conditions (4.1) and $|z|\leq|t-s|^{1/\alpha}$, using Corollary 4.4 and (3.21), we get
$\left|\delta_{\vartheta\sigma_{1}z}\nabla^{2}p^{(2),\sigma}_{s,t}(y)\right|$ $\lesssim|t-s|^{-1/2}|\vartheta\sigma_{1}z|\times|t-s|^{-\frac{2}{\alpha}}\varrho_{\alpha}(t-s,y)$
$\lesssim|t-s|^{-\frac{3}{\alpha}}|z|\varrho_{\alpha}(t-s,y),$